# Equivalent Circuit Modeling of Mutually Resistively Coupled Microwave Cavities with Enhanced Phase Sensitivity Using Thin Metallic Foils

### 05 Mar 2026 v1

@@ II Equivalent Circuit Model
Each of the parallel LCR circuits ($n=1,2,3$) has a self admittance of
@@
The mutual admittances $Y_{13}=Y_{31}=\frac{1}{R13}$ and $Y_{23}=Y_{32}=\frac{1}{R23}$ (simplification justified in Section IV) couple resonators 1–3 and 2–3, respectively, while resonators 1 and 2 remain uncoupled with no mutual foils ($Y_{12}=Y_{21}=0$). It is here that we can introduce a normalised mutual coupling term $\Delta$ with reference to the mutual resistance
$\Delta_{n3}=\frac{R_{03}\sqrt{\beta_{n}\beta_{3}}}{R_{n3}}.$ (6)
@@
$T=\frac{1}{\sqrt{2}}\frac{f_{3}\sqrt{\beta_{3}}(f_{1}\sqrt{\beta_{1}}\gamma_{2}\Delta_{13}+e^{i\phi}\alpha f_{2}\sqrt{\beta_{2}}\gamma_{1}\Delta_{23})}{-\gamma_{1}\gamma_{2}\gamma_{3}+f_{1}\gamma_{2}f_{3}\Delta_{13}^{2}+\gamma_{1}f_{2}f_{3}\Delta_{23}^{2}},$ (8)
@@ IV.1 Derivation
The dissipated power $Pd$, is generally
@@
$R_{n3}=\frac{1\pm\sqrt{1+4Y_{n3}^{2}R_{n}R_{3}}}{2Y_{n3}}\approx\frac{1}{Y_{n3}},$ (27)
with the approximation valid for $4Y_{n3}^{2}R_{n}R_{3}<<1$, indeed the case for the parameters shown in Table 2 and with frequencies around $6$ GHz. $R_{n}$ and $R_{3}$ are the parallel resistance of the respective resonator given by
@@ IV.2 Calculated Values
The mutual resistance $R_{n3}$ is calculated to lie in the range $R_{n3}\approx(1\text{--}51)\times 10^{11}\,\Omega$. The positive root of Eq. 27 is taken as the physical solution; at very small coupling the analytical expression may yield small negative values, so the magnitude is reported. This extremely large value is intuitively consistent with the parallel-circuit picture: only small current flows across the foil, so the resulting coupling is weak, when the foil thickness spans multiple skin depths. Recalling Eq. 6, and expressing the resonator resistance as $R_{n}=R_{0n}\beta_{n}$, the mutual coupling is calculated to lie in the range $\Delta_{n3}\approx(1\text{--}48)\times 10^{-6}$ when the estimated foil thickness of $(9\pm 1)\,\mu\mathrm{m}$ is considered. Comparing this with the experimentally fitted values of $\Delta_{13}=(5.00\pm 0.01)\times 10^{-6}$ and $\Delta_{23}=(4.10\pm 0.01)\times 10^{-6}$ shows good agreement within the expected uncertainty. Given the strong sensitivity of the coupling to foil thickness and other geometric parameters, agreement at this level constitutes successful verification of the model.

### 09 Apr 2026 v2

# Equivalent Circuit Modeling of Foil-Mediated Dissipative Coupling in Microwave Cavities with Enhanced Phase Response

@@ II Equivalent Circuit Model
Each of the parallel LCR circuits ($n=1,2,3$) has a self admittance approximated by $Y_{nn}\approx\frac{1}{Z_{nn}}$ in the small mutual coupling limit, where $Z_{nn}$ is the self impedance of cavity n. The admittance is given by
@@
The mutual admittances $Y_{13}=Y_{31}\approx\frac{-R_{13}}{R_{1}R_{3}}$ and $Y_{23}=Y_{32}\approx\frac{-R_{23}}{R_{2}R_{3}}$ (simplification justified in Section IV) couple resonators 1–3 and 2–3, respectively, while resonators 1 and 2 remain uncoupled with no mutual foils ($Y_{12}=Y_{21}=0$). It is here that we can introduce a normalised mutual coupling term $\Delta$ with reference to the mutual resistance
$\Delta_{n3}=\frac{R_{n3}}{R_{03}\sqrt{\beta_{n}\beta_{3}}}.$ (6)
@@
$T=-\frac{1}{\sqrt{2}}\frac{f_{3}\sqrt{\beta_{3}}(f_{1}\sqrt{\beta_{1}}\gamma_{2}\Delta_{13}+e^{i\phi}\alpha f_{2}\sqrt{\beta_{2}}\gamma_{1}\Delta_{23})}{-\gamma_{1}\gamma_{2}\gamma_{3}+f_{1}\gamma_{2}f_{3}\Delta_{13}^{2}+\gamma_{1}f_{2}f_{3}\Delta_{23}^{2}},$ (8)
@@ IV.1 Derivation
The dissipated power PdPd, is generally
@@
$R_{n3}=\frac{1\pm\sqrt{1+4Y_{n3}^{2}R_{n}R_{3}}}{2Y_{n3}}\approx-Y_{n3}R_{n}R_{3},$ (27)
with the approximation valid for $4Y_{n3}^{2}R_{n}R_{3}<<1$ in the negative root, indeed the case for the parameters shown in Table 2 and with frequencies around 66 GHz. $R_{n}$ and $R_{3}$ are the parallel resistance of the respective resonator given by
@@ IV.2 Calculated Values
The mutual resistance $R_{n3}$ is calculated to lie in the range $R_{n3}\approx(6\text{--}300)\,\Omega$. The negative root of Eq. 27 is taken as the physical solution; the sign not meaningful, so the magnitude is reported. This extremely small value is intuitively consistent with the parallel-circuit picture: only small current flows across the foil, so the resulting coupling is weak, when the foil thickness spans multiple skin depths. Recalling Eq. 6, and expressing the resonator resistance as $R_{n}=R_{0n}\beta_{n}$, the mutual coupling is calculated to lie in the range $\Delta_{n3}\approx(1\text{--}48)\times 10^{-6}$ when the estimated foil thickness of $(9\pm 1)\,\mu\mathrm{m}$ is considered. Comparing this with the experimentally fitted values of $\Delta_{13}=(5.00\pm 0.01)\times 10^{-6}$ and $\Delta_{23}=(4.10\pm 0.01)\times 10^{-6}$ shows good agreement within the expected uncertainty. Given the strong sensitivity of the coupling to foil thickness and other geometric parameters, agreement at this level constitutes successful verification of the model.